# Groups of matrices with approximately submultiplicative spectra

### v3

@@ 1. Introduction
A semigroup $\mathcal{S}$ of complex $n\times n$ matrices is said to have submultiplicative spectrum if for every pair $A,B\in\mathcal{S}$, every eigenvalue of the product $AB$ is equal to a product of an eigenvalue of $A$ and an eigenvalue of $B$. This property was introduced by Lambrou, Longstaff, and Radjavi in 1992 [8] and has since been extensively studied. It has led to numerous nice structure results for matrix groups and semigroups. Irreducible semigroups with this property are essentially finite nilpotent groups [12, Thm. 3.3.4, Thm. 3.3.5] (see also [8]). In the original paper [8] the authors proved that such irreducible groups exist in all odd dimensions. Kramar considered even dimensions in [4] and showed that for even $n$, irreducible groups of $n\times n$ matrices with submultiplicative spectra exist if and only if $n$ is divisible by $8$. In [11] the structure of irreducible $2$-groups was studied and in [5] this study was extended to include $p$-groups for general $p$. In [6] the class ($\hat{\mathrm{s}}$) of groups $G$ with the property that all their irreducible sub-representations are submultiplicative was introduced. A systematic study of this class of finite groups was initiated in [3].
In this note we start the study of an approximate version of submultiplicativity. The idea of replacing exact conditions with approximate ones has a rich (recent) history. The first “approximate version” result in the context of simultaneous triangularization of matrix semigroups is [2]. There the authors ask how small can the spectra of nonzero commutators in a unitary group of matrices be? They show that for a non-commutative unitary matrix group $\mathcal{G}$, there always exist elements $A,B\in\mathcal{G}$ such that the spectral radius $\rho(AB-BA)$ of their ring commutator is at least $\sqrt{3}$. Alternatively, this means that for unitary groups, the exact triangularizing condition (actually diagonalizing in this case)
@@
for some $\varepsilon_{f,n}>0$ that depends on $f$ and $n$. Naturally in concrete cases, it can also be important to find the largest possible $\varepsilon_{f,n}$. In particular, the question of whether $\varepsilon_{f,n}$ is bounded below by some polynomial in $\frac{1}{n}$ is of interest.
We say that a semigroup $\mathcal{S}$ of complex $n\times n$ matrices is $\varepsilon$-submuliplicative if for all $A,B\in\mathcal{S}$ and for every eigenvalue $\gamma$ of $AB$, there exists eigenvalues $\alpha$ of $A$ and $\beta$ of $B$ such that
@@
It is almost immediate that irreducible such groups do not contain any nonzero nilpotent elements. However, they need not be groups. For any value of $\varepsilon$, there are irreducible rank-one semigroups that satisfy this condition. In this note we focus primarily on unitary groups. In this case the condition becomes
@@
(Here $\arg(z)\in(-\pi,\pi]$ denotes the principal argument of $z$.) We show, among other things, that for a unitary group, and $\varepsilon=\frac{1}{2n^{2}}$, the $\varepsilon$-argument-submultiplicativity implies that the group is finite modulo its centre. We show, by example, that the order of $\varepsilon$ above is sharp: there exist $\frac{2}{n^{2}}$-ASM groups of unitary matrices that are not finite modulo their centre. We do not know at this point whether the quadratic order is still sharp if we additionally assume irreducibility. More precisely: there are easy examples of irreducible $\varepsilon$-ASM groups that are not essentially finite for $\varepsilon>\frac{1}{2n}$, but we do not know whether such examples exist when $\varepsilon=\frac{c}{n^{2}}$ for some fixed constant $c$.
Our approximate results are quite different from those in literature. Most similar sounding results deal with studying conditions that imply reducibility or triangularizability of matrix collections (groups, semigroups, etc.). In the case of submultiplicativity (and its approximate version) the situation is different. We mainly deal with irreducible (or at least completely reducible) groups and semigroups; the condition in question then implies something about their structure. Another important distinction is that in most other situations in the literature the approximate condition actually implies the exact condition for small $\varepsilon$ (and not much is said about the structure when $\varepsilon$ is not sufficiently small to ensure this). In our paper, this is not the case. Even for an $\varepsilon$ that is sufficiently small to imply structure results (e.g., essential finiteness) we have interesting examples of matrix groups that are approximately submultiplicative, but not submultiplicative.
@@ 2. Preliminaries
Define semigroup $\mathcal{S}_{r}$ of matrices for $10<r<$ 1 as
@@
It has been shown in [2, Ex. 2.3.] that $\mathcal{S}_{r}$ is irreducible. It is easy to see that $\mathcal{S}_{r}$ cannot be essentially finite. Indeed, it is obvious that two distinct elements of $\mathcal{S}_{r}$ whose $(1,1)$-entries are both equal to $1$ cannot be multiples of each other and (again obviously) there are infinitely many pairwise-distinct such elements. It is fairly straightforward that $\mathcal{S}_{r}$ is $\frac{4r^{2}}{(1-r^{2})^{2}}$-submultiplicative (see below). Note that for any $\varepsilon>0$ we have that for a sufficiently small $r$, $\frac{4r^{2}}{(1-r^{2})^{2}}<\varepsilon.$ Let
@@
Then the unique nonzero eigenvalue of $A$ is $\alpha=\lambda(1+x^{*}y)$, the unique nonzero eigenvalue of $B$ is $\beta=\mu(1+a^{*}b)$, and the unique nonzero eigenvalue of $AB$ is $\gamma=\lambda\mu(1+x^{*}b)(1+a^{*}y)$. A quick direct computation shows that
@@
$=$ $\left|\frac{(1+x^{*}b)(1+a^{*}y)}{(1+x^{*}y)(1+a^{*}b)}-1\right|\leq\left|\left|\frac{(1+x^{*}b)(1+a^{*}y)}{(1+x^{*}y)(1+a^{*}b)}\right|-1\right|.$
Note that each of the factors $|(1+a^{*}b)|,|(1+a^{*}y)|,|(1+x^{*}b)|,|(1+x^{*}y)|$ lies in the interval $\left[1-r^{2},1+r^{2}\right]$ and hence the value of $\left|\frac{(1+x^{*}b)(1+a^{*}y)}{(1+x^{*}y)(1+a^{*}b)}\right|$ must be between $\frac{(1-r^{2})^{2}}{(1+r^{2})^{2}}$ and $\frac{(1+r^{2})^{2}}{(1-r^{2})^{2}}$. Subtracting $1$ from these values then yields the promised estimate.
@@ 3. Main Results
As mentioned in the introduction, a scaled multiplicative version of approximate submultiplicativity is more convenient in the context of unitary groups.
@@
We say that the spectrum of a group $\mathcal{G}\subseteq\mathcal{M}_{n}(\mathbb{C})$ of unitary matrices is $\varepsilon^{\prime}$-argument-submultiplicative (or $\varepsilon^{\prime}$-ASM) if for every $A,B\in\mathcal{G}$ and every $\gamma\in\sigma(AB)$ there exists an $\alpha\in\sigma(A)$ and a $\beta\in\sigma(B)$ such that
@@
But first we show by example, that for infinitely many $n$, there exists $\frac{2}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that are not finite modulo their centres. Let $p$ be an odd prime, let $D$ be any $p\times p$ unitary diagonal matrix of determinant $1$, and let $C$ to be the $p\times p$ cycle matrix, i.e.,
@@
We will use the following well known fact about the spectra of $DC^{k}$. We include a sketch of the proof for the sake of completeness.
@@
Alternatively, it is in fact possible to see directly, that matrices $DC^{k}$ and $C^{k}$ are similar: the matrix form of the linear map corresponding to $DC^{k}$ in basis $f_{j}=\left(\prod_{i=1}^{j}d_{i}\right)e_{j}$, $j=1,\ldots,p$ is $C^{k}$ (here $d_{i}$ denotes the $i$-th diagonal entry of $D$). It is also well known that matrices $C^{k}$ and $C$ are similar by the change of basis $g_{j}=e_{1+(j-1)k}$, $j=1,\ldots,p$ (using the convention that for $j>p$ we have $e_{j}=e_{j-p}$). ∎
@@
Let (as above) $C$ be the $p\times p$ cycle matrix and $D$ a unitary diagonal matrix of determinant one. Furthermore, let $\xi=e^{\frac{2\pi i}{p^{2}}}$ (i.e., $\xi$ is a fixed primitive $p^{2}$-th root of unity), $k\in\{0,1,2,\dots,p-1\}$, and $a_{j}\in\{1,2,\dots,p-1\}$ for $j=1,2,\dots,p-1$. For any choice of $k$, $D$, $(a_{j})_{j=1}^{p-1}$ we define the corresponding “tadopole matrix” $A\in\mathcal{M}_{2p}(\mathbb{C})$ by
@@
to denote the “tail” of the tadpole matrix. With this notation we write (purposefully abusing the notation somewhat):
@@
Let $\mathcal{T}$ be the set of all such tadpole matrices (i.e., tadpole matrices corresponding to all possible choices of $D$, $k$, $(a_{j})_{j=1}^{p-1}$). We claim that that $\mathcal{T}$ forms a group under matrix multiplication. Clearly the identity matrix is in $\mathcal{T}$ and inverses of elements in $\mathcal{T}$ are again in $\mathcal{T}$. We are left to show that a product of any two elements of $\mathcal{T}$ is again in $\mathcal{T}$. Let $A=A\left(D,k,(a_{j})_{j=1}^{p-1}\right)$ be as above, and let $B\in\mathcal{T}$ be written as
@@
Case 1. Suppose both $A$ and $B$ are diagonal. In this case we obviously have $\sigma(AB)\subseteq\sigma(A)\sigma(B)$.
@@
The tail of $AB$ is the product of the tails of $A$ and $B$. Since these are diagonal matrices, we have that $\sigma(A_{T}B_{T})\subseteq\sigma(A_{T})\sigma(B_{T})$. We are left to prove the same for the heads of $A$ and $B$. Since $D_{A}$ and $D_{B}$ are unitary, diagonal matrices of determinant $1$, $D_{A}D_{B}$ is also a unitary diagonal matrix of determinant $1$. By Lemma 3 we have that
@@
Case 4. Suppose that neither $A$ nor $B$ be is diagonal, but $AB$ is. So $k\neq 0$, $\ell=p-k$ and $r=0$. We will first show that $\sigma(A)\sigma(B)=\{1,\xi,\xi^{2},\dots,\xi^{p^{2}-1}\}$. Since neither $A$ nor $B$ is diagonal, the spectra of their heads consist of all $p$-th roots of unity. The spectra of tails are contained in the set of all $p^{2}$-th roots of unity and hence $\sigma(A)\sigma(B)$ must be a subset of the set of all powers of $\xi$. We are left to show that $\sigma(A)\sigma(B)$ contains $\{1,\xi,\xi^{2},\dots,\xi^{p^{2}-1}\}$. We know from the above observations that:
$\sigma(A)\supseteq\{1,\xi^{k+a_{1}p},\xi^{2k+a_{2}p},\dots,\xi^{(p-1)k+a_{p-1}p}\}$
$\sigma(B)\supseteq\{1,\xi^{(-a_{j}+t)p}:j=1,2,\dots,p-1,t=0,1,\dots,p-1\}$
@@
(Take $\alpha,\beta$ to be such that $\alpha\beta$ is the $p^{2}$-th root of $1$ that is closest to $\gamma$.) ∎
The theorem above shows that for $n=2p$ where $p$ is an odd prime, there exists a group $\mathcal{G}$ of unitary $n\times n$ matrices that is $\frac{2}{n^{2}}$-ASM, but not finite modulo its centre. In the next result we show that reducing the size of $\varepsilon^{\prime}=\frac{2}{n^{2}}$ by a factor of $4$ does in fact force the group in question to be finite modulo its centre.
@@
Let $\mathcal{G}\subseteq\mathcal{M}_{n}(\mathbb{C})$ be a group of unitary matrices. If $\mathcal{G}$ is $\frac{1}{2n^{2}}$-argument-submultiplicative then $\mathcal{G}$ must be finite modulo its centre.
@@ 4.1. Optimality of our results
The group of tadpole matrices in Example 4 is an example of an $\frac{2}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that is not finite modulo its centre (here $n=2p$, where $p$ is an odd prime). But this group is clearly not irreducible. The group of its heads (i.e., the group of $p\times p$ matrices generated by all unitary diagonal matrices $D$ and the cycle matrix $C$) is easily seen to be an example of an irreducible group of unitary $p\times p$-matrices that is $\frac{1}{2p}$-ASM, but not essentially finite. Can we do better? In particular: is there some constant $c$ such that for infinitely many $n$, there exists irreducible $\frac{c}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that are not essentially finite?

### v4

@@ 1. Introduction
A semigroup $\mathcal{S}$ of complex $n\times n$ matrices has a submultiplicative spectrum if for every pair $A,B\in\mathcal{S}$, every eigenvalue of the product $AB$ is equal to a product of an eigenvalue of $A$ and an eigenvalue of $B$. This property was introduced by Lambrou, Longstaff, and Radjavi in 1992 [8] and has since been extensively studied. It has led to numerous nice structure results for matrix groups and semigroups. Irreducible semigroups with this property are essentially finite nilpotent groups [12, Thm. 3.3.4, Thm. 3.3.5] (see also [8]). In the original paper [8], the authors proved that such irreducible groups exist in all odd dimensions. Kramar considered even dimensions in [4] and showed that for even $n$, there exist irreducible groups of $n\times n$ matrices with submultiplicative spectra if and only if $n$ is divisible by $8$. In [11] the structure of irreducible $2$-groups was studied and in [5] this study was extended to include $p$-groups for general $p$. In [6] the class ($\hat{\mathrm{s}}$) of groups $G$ with the property that all their irreducible sub-representations are submultiplicative was introduced. A systematic study of this class of finite groups was initiated in [3].
In this note, we start the study of an approximate version of submultiplicativity. The idea of replacing exact conditions with approximate ones has a rich (recent) history. The first “approximate version” result in the context of simultaneous triangularization of matrix semigroups is [2]. There, the authors ask “how small can the spectra of nonzero commutators in a unitary group of matrices be?” They show that for a non-commutative unitary matrix group $\mathcal{G}$, there always exist elements $A,B\in\mathcal{G}$ such that the spectral radius $\rho(AB-BA)$ of their ring commutator is at least $\sqrt{3}$. Alternatively, this means that for unitary groups, the exact triangularizing condition (actually diagonalizing in this case)
@@
for some $\varepsilon_{f,n}>0$ that depends on $f$ and $n$. Naturally, in concrete cases, it can also be important to find the largest possible $\varepsilon_{f,n}$. In particular, the question of whether $\varepsilon_{f,n}$ is bounded below by some polynomial in $\frac{1}{n}$ is of interest.
We say that a semigroup $\mathcal{S}$ of complex $n\times n$ matrices is $\varepsilon$-submuliplicative if for all $A,B\in\mathcal{S}$ and for every eigenvalue $\gamma$ of $AB$, there exist eigenvalues $\alpha$ of $A$ and $\beta$ of $B$ such that
@@
It is almost immediate that irreducible such semigroups do not contain any nonzero nilpotent elements. However, such semigroups need not be groups. For any value of $\varepsilon>0$, there are irreducible semigroups of matrices of rank at most $1$, that satisfy this condition. In this note we focus primarily on unitary groups. In this case, the condition becomes
@@
(Here $\arg(z)\in(-\pi,\pi]$ denotes the principal argument of $z$.) We show, among other things, that for a unitary group, and $\varepsilon=\frac{1}{2n^{2}}$, the $\varepsilon$-argument-submultiplicativity implies that the group is finite modulo its centre. We show, by example, that the order of $\varepsilon$ above is sharp: there exist $\frac{2}{n^{2}}$-ASM groups of unitary matrices that are not finite modulo their centre. We do not know at this point whether the quadratic order is still sharp if we additionally assume irreducibility. More precisely: there are easy examples of irreducible $\varepsilon$-ASM groups that are not essentially finite for $\varepsilon>\frac{1}{2n}$, but we do not know if such examples exist when $\varepsilon=\frac{c}{n^{2}}$ for some fixed constant $c$.
Our approximate results are quite different from those in the literature. Most similar sounding results deal with studying conditions that imply reducibility or triangularizability of matrix collections (groups, semigroups, etc.). In the case of submultiplicativity (and its approximate version) the situation is different. We mainly deal with irreducible (or at least completely reducible) groups and semigroups; the condition in question then implies something about their structure. Another important distinction is that in most other situations in the literature, the approximate condition actually implies the exact condition for small $\varepsilon$ (and not much is said about the structure when $\varepsilon$ is not sufficiently small to ensure this). In our paper, this is not the case. Even for an $\varepsilon$ that is sufficiently small to imply structure results (e.g., essential finiteness) we have interesting examples of matrix groups that are approximately submultiplicative, but not submultiplicative.
@@ 2. Preliminaries
Define the semigroup $\mathcal{S}_{r}$ of matrices for $10<r<$ 1 as
@@
In [2, Ex. 2.3.] it has been shown that $\mathcal{S}_{r}$ is irreducible. It is easy to see that $\mathcal{S}_{r}$ cannot be essentially finite. Indeed, it is obvious that two distinct elements of $\mathcal{S}_{r}$ whose $(1,1)$-entries are both equal to $1$ cannot be multiples of each other and (again obviously) there are infinitely many pairwise-distinct such elements. It is fairly straightforward that $\mathcal{S}_{r}$ is $\frac{4r^{2}}{(1-r^{2})^{2}}$-submultiplicative (see below). Note that for any $\varepsilon>0$ we have that for a sufficiently small $r$, $\frac{4r^{2}}{(1-r^{2})^{2}}<\varepsilon.$ Let
@@
Then the unique nonzero eigenvalue of $A$ is $\alpha=\lambda(1+a^{*}b)$, the unique nonzero eigenvalue of $B$ is $\beta=\mu(1+x^{*}y)$, and the unique nonzero eigenvalue of $AB$ is $\gamma=\lambda\mu(1+a^{*}y)(1+x^{*}b)$. A quick direct computation shows that
@@
$=$ $\left|\frac{(1+a^{*}y)(1+x^{*}b)}{(1+a^{*}b)(1+x^{*}y)}-1\right|\leq\left|\left|\frac{(1+a^{*}y)(1+x^{*}b)}{(1+a^{*}b)(1+x^{*}y)}\right|-1\right|.$
Note that each of the factors $|(1+a^{*}b)|,|(1+a^{*}y)|,|(1+x^{*}b)|,|(1+x^{*}y)|$ lies in the interval $\left[1-r^{2},1+r^{2}\right]$ and therefore the value of $\left|\frac{(1+x^{*}b)(1+a^{*}y)}{(1+x^{*}y)(1+a^{*}b)}\right|$ must be between $\frac{(1-r^{2})^{2}}{(1+r^{2})^{2}}$ and $\frac{(1+r^{2})^{2}}{(1-r^{2})^{2}}$. Subtracting $1$ from these values then yields the promised estimate.
@@ 3. Main Results
As mentioned in the introduction, a scaled multiplicative version of approximate submultiplicativity is more convenient in the context of unitary groups: in this case the eigenvalues always lie on the unit circle in $\mathbb{C}$ and we find it more convenient to consider the “arc-distance” between them. The scaling factor $\frac{1}{2\pi}$ was chosen so that the distance between two consecutive $n$-th roots of unity is $\frac{1}{n}$ (and, more generally, for $x\in[0,1)$, the distance between $1$ and $e^{2\pi xi}$ is $x$).
@@
We say that the spectrum of a group $\mathcal{G}\subseteq\mathcal{M}_{n}(\mathbb{C})$ of unitary matrices is $\varepsilon^{\prime}$-argument-submultiplicative (or $\varepsilon^{\prime}$-ASM) if for every $A,B\in\mathcal{G}$ and every $\gamma\in\sigma(AB)$, there exist elements $\alpha\in\sigma(A),\beta\in\sigma(B)$ such that
@@
But first we show by example, that for infinitely many $n$, there exist $\frac{2}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that are not finite modulo their centres. Let $p$ be an odd prime, let $D$ be any $p\times p$ unitary diagonal matrix of determinant $1$, and let $C$ to be the $p\times p$ cycle matrix, i.e.,
@@
We will use the following well-known fact about the spectra of $DC^{k}$. We include a sketch of the proof for the sake of completeness.
@@
Alternatively, it is in fact possible to see directly, that the matrices $DC^{k}$ and $C^{k}$ are similar: the matrix form of the linear map corresponding to $DC^{k}$ in basis $f_{j}=\left(\prod_{i=1}^{j}d_{i}\right)e_{j}$, $j=1,\ldots,p$ is $C^{k}$ (here $d_{i}$ denotes the $i$-th diagonal entry of $D$). It is also well-known that the matrices $C^{k}$ and $C$ are similar by the change of basis $g_{j}=e_{1+(j-1)k}$, $j=1,\ldots,p$ (here $(e_{j})_{j=1}^{p}$ denotes the standard orthonormal basis of $\mathbb{C}^{p}$, we use the convention that for $j>p$ we have $e_{j}=e_{j-p}$). ∎
@@
Let (as above) $C$ be the $p\times p$ cycle matrix and let $\xi=e^{\frac{2\pi i}{p^{2}}}$ (i.e., $\xi$ is a fixed primitive $p^{2}$-th root of unity). For a unitary diagonal matrix $D$, and integer parameters $k\in\{0,1,2,\dots,p-1\}$, and $a_{j}\in\{0,1,2,\dots,p-1\}$ with $j=1,2,\dots,p-1$, we define the corresponding “tadopole matrix” $A\in\mathcal{M}_{2p}(\mathbb{C})$ by
@@
to denote the “tail” of the tadpole matrix. With this notation in mind, we write (purposefully abusing the notation somewhat):
@@
Let $\mathcal{T}$ be the set of all such tadpole matrices (i.e., tadpole matrices corresponding to all possible choices of $D$, $k$, $(a_{j})_{j=1}^{p-1}$). We claim that $\mathcal{T}$ forms a group under matrix multiplication. Clearly, the identity matrix is in $\mathcal{T}$ and the inverses of elements from $\mathcal{T}$ are again in $\mathcal{T}$. We are left to show that the product of any two elements of $\mathcal{T}$ is also in $\mathcal{T}$. Let $A=A\left(D,k,(a_{j})_{j=1}^{p-1}\right)$ be as above and let $B\in\mathcal{T}$ be written as
@@
Case 1. Suppose that both $A$ and $B$ are diagonal. In this case we obviously have $\sigma(AB)\subseteq\sigma(A)\sigma(B)$.
@@
The tail of $AB$ is the product of the tails of $A$ and $B$. Since these are diagonal matrices, we have that $\sigma(A_{T}B_{T})\subseteq\sigma(A_{T})\sigma(B_{T})$. We are left to prove the same for the heads of $A$ and $B$. Since $D_{A}$ and $D_{B}$ are unitary diagonal matrices of determinant $1$, $D_{A}D_{B}$ is also a unitary diagonal matrix of determinant $1$. By Lemma 3 we have that
@@
Case 4. Suppose that neither $A$ nor $B$ is diagonal, but $AB$ is. So $k\neq 0$, $\ell=p-k$ and $r=0$. We will first show that $\sigma(A)\sigma(B)=\{1,\xi,\xi^{2},\dots,\xi^{p^{2}-1}\}$. Since neither $A$ nor $B$ is diagonal, the spectra of their heads consist of all $p$-th roots of unity. The spectra of tails are contained in the set of all $p^{2}$-th roots of unity and hence $\sigma(A)\sigma(B)$ must be a subset of the set of all powers of $\xi$. We are left to show that $\sigma(A)\sigma(B)$ contains $\{1,\xi,\xi^{2},\dots,\xi^{p^{2}-1}\}$. We know from the above observations that:
$\sigma(A)$ $\supseteq$ $\sigma(A_{T})=\{1,\xi^{k+a_{1}p},\xi^{2k+a_{2}p},\dots,\xi^{(p-1)k+a_{p-1}p}\},$
$\sigma(B)$ $\supseteq$ $\sigma(B_{H})=\{1,\theta,\ldots,\theta^{p-1}\}=\{\xi^{tp}:t=0,1,\ldots,p-1\}$
$\supseteq$ $\{1,\xi^{(-a_{j}+t)p}:j=1,2,\dots,p-1,t=0,1,\dots,p-1\}.$
@@
(Take $\alpha,\beta$ to be such that the product $\alpha\beta$ is the $p^{2}$-th root of $1$ that is closest to $\gamma$.) ∎
The above theorem shows that for $n=2p$ where $p$ is an odd prime, there exists a group $\mathcal{G}$ of unitary $n\times n$ matrices that is $\frac{2}{n^{2}}$-ASM, but not finite modulo its centre. In the next result, we show that reducing the size of $\varepsilon^{\prime}=\frac{2}{n^{2}}$ by a factor of $4$ does in fact force the group in question to be finite modulo its centre.
@@
Let $\mathcal{G}\subseteq\mathcal{M}_{n}(\mathbb{C})$ be a group of unitary matrices. If $\mathcal{G}$ is $\frac{1}{2n^{2}}$-argument-submultiplicative, then $\mathcal{G}$ must be finite modulo its centre.
@@ 4.1. Optimality of our results
The group of tadpole matrices in Example 4 is an example of a $\frac{2}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that is not finite modulo its centre (here $n=2p$, where $p$ is an odd prime). But this group is clearly not irreducible. The group of its heads (i.e., the group of $p\times p$ matrices generated by all unitary diagonal matrices $D$ and the cycle matrix $C$) is easily seen to be an example of an irreducible group of unitary $p\times p$-matrices that is $\frac{1}{2p}$-ASM, but not essentially finite. Can we do better? In particular: is there some constant $c$ such that for infinitely many $n$, there exist irreducible $\frac{c}{n^{2}}$-ASM groups of unitary $n\times n$ matrices that are not essentially finite?